A load of mass 150kg is to be pulled up a smooth plane at 15 degrees to the vertical. Calculate the minimum force required to pull it up the plane if the plane is 75%efficient


1
Expert's answer
2021-03-29T09:01:20-0400

Solution.

"m=150kg;"

"\alpha=15^o;"

"\eta=75%;" %;

"F-?;"

"\eta=\dfrac{A}{A_f}\sdot100" % "\implies A=\dfrac{\eta A_f}{100\%};"

"A=mgh;"

"A_f=Fl;"

"sin(90-\alpha)=\dfrac{h}{l} \implies h=lsin(90-\alpha);"

"mglsin(90-\alpha)=\dfrac{\eta Fl}{100\%};"

"mgsin(90-\alpha)=\dfrac{\eta F}{100\%}\implies F=\dfrac{mgsin(90-\alpha)\sdot100\%}{\eta};"

"F=\dfrac{150kg\sdot9.8ms^{-2}\sdot 0.966\sdot100\%}{75\%}=1893.36N;"


Answer:"F=1893.36N."